# Energy-Aware Lane Planning for Connected Electric Vehicles in Urban Traffic: Design and Vehicle-in-the-Loop Validation

### v3

@@ V-A3 Constraints
Collision avoidance constraints with respect to the NPC vehicles are also considered, and their set representation is denoted as $\mathcal{F}^{LK}(s^{\text{front}},v^{\text{front}})$, where $s^{\text{front}}$ and $v^{\text{front}}$ denote the states of the vehicle immediately preceding the ego vehicle in the same lane. We assume that the NPC vehicles maintain constant velocity during the prediction horizon $N$. For more details, refer to [arpae-lc].
@@
$\mathcal{P}(s_{tl}^{\text{next}})=\{x\in\mathbb{R}^{2}|s\leq s_{tl}^{\text{next}}+d\},$ (14)
$\mathcal{S}(s_{tl}^{\text{next}})=\{x\in\mathbb{R}^{2}|s_{tl}^{\text{next}}-d\leq s\},$ (15)
@@
where $\bar{t}_{\text{pass}}$ and $\underline{t}_{\text{pass}}$ denote the upper and lower bounds of the time interval during which the pass constraint is active; analogous definitions apply to the stop constraint. The parameter $t_{g}^{\text{margin}}$ is a user-defined safety margin that ensures the vehicle passes through the traffic light within the green phase. In this work, we set $t_{g}^{\text{margin}}=2$ s. Additionally, the function $CT$ denotes the cycle time adjustment, which accounts for the current signal phase of the upcoming traffic light and the decision to either pass or stop. Let $\phi\in\{0,1,2\}$ be the categorical encoding of the next traffic light’s signal phase, corresponding to green, yellow, and red respectively: $0\mapsto g,\ 1\mapsto y,\ 2\mapsto r]$. The next traffic light’s cycle durations is $\text{Cycle}:=[t^{\text{next}}_{g},t^{\text{next}}_{y},t^{\text{next}}_{r}],$ where $t^{\text{next}}_{g}$, $t^{\text{next}}_{y}$, and $t^{\text{next}}_{r}$ denote the durations of the green, yellow, and red phases respectively—which can be obtained from the traffic light’s SPaT data. The $CT$ function is defined as follows:
@@ V-A4 Optimal Control Problem
where $x_{k|t}$ denotes the decision variable at time step $k$ predicted at time $t$ and $N$ is the prediction horizon. In this work, we use $N=140$ to achieve 1 $Hz$ lane selection planning with the available hardware. The $\mathbb{I}_{k_{1}}^{k_{2}}$ notation denotes the set $\{k_{1},k_{1}+1,\dots,k_{2}\}$. The optimal control problem has three components for the cost function. The first component is the energy minimization cost which uses the data-driven energy consumption model: $\mathcal{E}$, which is quadratic and convex w.r.t. the decision variables. The second component $J^{LK}_{\text{smooth}}(\{x_{k|t}\}^{t+N-1}_{t})$ is the stage cost added for smoothness of the longitudinal trajectory and passenger comfort [arpae-lc]. After the optimal control problem is formulated in CasADi [casadi] and solved using IPOPT [ipopt], the passing velocity of the first traffic light $v^{\star}_{\text{pass}}$ is extracted from the optimal solution, which is used to initialize the graph (9). After computing (11), we define
@@ VI-2 Experiment Results
As a result, the proposed algorithm achieved a 39% reduction in motion energy consumption compared to Baseline 1, averaging $580$ Wh versus $808$ Wh. Furthermore, it achieved a 24% reduction in total energy consumption (including auxiliary power usage) compared to Baseline 1 on average. While one might argue that the energy savings are simply due to the vehicle driving more slowly, the proposed algorithm actively computed energy-efficient velocity profiles and lane selections based on an energy cost function. No lower bound was imposed on the velocity profile, allowing the algorithm to explore the energy-optimal behavior of the electric vehicle. Although Baseline 2 exhibited similar motion energy consumption to the proposed algorithm (within 2% on average), its significantly longer trip time led to higher overall energy usage due to increased auxiliary power consumption, resulting in only a 7% reduction on average. Ultimately, the proposed algorithm demonstrated the lowest total energy consumption as defined by the cost function.

### v4

@@ V-A3 Constraints
Collision avoidance constraints with respect to the NPC vehicles are also considered, and their set representation is denoted as $\mathcal{F}^{LK}(s^{\text{front}},v^{\text{front}})$, where $s^{\text{front}}$ and $v^{\text{front}}$ denote the states of the vehicle immediately preceding the ego vehicle in the same lane. We assume that the NPC vehicles maintain constant velocity during the prediction horizon NN. For more details, refer to [arpae-lc].
@@
$\mathcal{P}(s_{tl}^{\text{next}})=\{x\in\mathbb{R}^{2}|s\geq s_{tl}^{\text{next}}+d\},$ (14)
$\mathcal{S}(s_{tl}^{\text{next}})=\{x\in\mathbb{R}^{2}|s_{tl}^{\text{next}}-d\geq s\},$ (15)
@@
where $\bar{t}_{\text{pass}}$ and $\underline{t}_{\text{pass}}$ denote the upper and lower bounds of the time interval during which the pass constraint is active; analogous definitions apply to the stop constraint. The parameter $t_{g}^{\text{margin}}$ is a user-defined safety margin that ensures the vehicle passes through the traffic light within the green phase. In this work, we set $t_{g}^{\text{margin}}=2$ s. Additionally, the function CTCT denotes the cycle time adjustment, which accounts for the current signal phase of the upcoming traffic light and the decision to either pass or stop. Let $\phi\in\{0,1,2\}$ be the categorical encoding of the next traffic light’s signal phase, corresponding to green, yellow, and red respectively: $0\mapsto g,\ 1\mapsto y,\ 2\mapsto r]$. The next traffic light’s cycle durations is $\text{Cycle}:=[t^{\text{next}}_{g},t^{\text{next}}_{y},t^{\text{next}}_{r}],$ where $t^{\text{next}}_{g}$, $t^{\text{next}}_{y}$, and $t^{\text{next}}_{r}$ denote the durations of the green, yellow, and red phases respectively—which can be obtained from the traffic light’s SPaT data. The CTCT function is defined as follows:
@@ V-A4 Optimal Control Problem
where $x_{k|t}$ denotes the decision variable at time step $k$ predicted at time $t$ and NN is the prediction horizon. In this work, we use $N=140$ to achieve 1 HzHz lane selection planning with the available hardware. The $\mathbb{I}_{k_{1}}^{k_{2}}$ notation denotes the set $\{k_{1},k_{1}+1,\dots,k_{2}\}$. The optimal control problem has three components for the cost function. The first component is the energy minimization cost which uses the data-driven energy consumption model: $\mathcal{E}$, which is quadratic and convex w.r.t. the decision variables. The second component $J^{LK}_{\text{smooth}}(\{x_{k|t}\}^{t+N-1}_{t})$ is the stage cost added for smoothness of the longitudinal trajectory and passenger comfort [arpae-lc]. After the optimal control problem is formulated in CasADi [casadi] and solved using IPOPT [ipopt], the passing velocity of the first traffic light $v^{\star}_{\text{pass}}$ is extracted from the optimal solution, which is used to initialize the graph (9). After computing (11), we define
@@ VI-2 Experiment Results
As a result, the proposed algorithm achieved a 39% reduction in motion energy consumption compared to Baseline 1, averaging 580580 Wh versus 808808 Wh. Furthermore, it achieved a 24% reduction in total energy consumption (including auxiliary power usage) compared to Baseline 1 on average. While one might argue that the energy savings are simply due to the vehicle driving more slowly, the proposed algorithm actively computed energy-efficient velocity profiles and lane selections based on an energy cost function. No lower bound was imposed on the velocity profile, allowing the algorithm to explore the energy-optimal behavior of the electric vehicle. Although Baseline 2 exhibited similar motion energy consumption to the proposed algorithm (within 2% on average), its significantly longer trip time led to higher overall energy usage due to increased auxiliary power consumption, resulting in only a 7% reduction on average. Ultimately, the proposed algorithm demonstrated the lowest total energy consumption as defined by the cost function.